# On zero-shot learning in neural state estimation of power distribution systems

### v1

@@ Abstract
This paper addresses the challenge of neural state estimation (NSE) in power distribution systems. We identified a research gap in the current state of the art, which lies in the inability of models to adapt to changes in the power grid, such as loss of sensors and branch switching. Our experiments demonstrate that graph neural networks (GNNs) are the most promising class of models for this use case, but also that their performance can degrade with scale. We propose augmentations to remedy this issue and perform a comprehensive grid search of different model configurations for common zero-shot learning (ZSL) scenarios in NSE.
Keywords:
@@
neural state estimation zero-shot learning transfer learning graph neural networks
1 Introduction
@@
Power system state estimation (PSSE) is the task of inferring the “state” of an electrical power grid from real-time data collected by various sensors distributed across the system. The “state” in this context generally refers to the voltage magnitudes and phase angles at each bus in the grid. These estimates are utilized in many control center applications, such as contingency analysis, load flow analysis, and optimal power flow (OPF).
@@
For many years, PSSE was mainly performed for the transmission grids using simplifying assumptions such as near-DC power flow and computational methods with poor scalability. This is enabled by balanced operation with a relatively simple, predominantly linear topology of transmission grids, given their scale and structure. It is also economically feasible to equip them with sufficient sensors [16].
@@
On the contrary, distribution grids, which transport electricity from substations to end consumers, present distinct challenges. Their unbalanced nature, radial or weakly meshed topology, high R/X ratios, and the inability to achieve sufficient sensor coverage complicate the state estimation process. Initially designed with transmission systems in mind, conventional methods often struggle to provide accurate state estimation in these more complex, dynamic, and less predictable distribution systems [20].
@@
However, with the proliferation of distributed energy resources (DERs) and other complex consumers, grid operators are facing the necessity of performing PSSE for distribution grids. Additionally, §14a of the German Energy Industry Act [5] effectively requires operators to develop transparency in distribution grids in order to align consumption with production from renewable energy sources, which requires PSSE.
Another example where a high-quality PSSE is required is model-based deep reinforcement learning (DRL). DRL has, in the past, already been successfully employed to aid in resilient power grid operation. Here, a world model is necessary that delivers precise results for partial observability environments: It speeds up training by allowing the agent to optimize its strategy based on an internal representation; in an actual real-world deployment, the agent not only profits from the full-observabilty the world model can provide, but it also allows to check for premature errors that could arise from DRL policy errors [18]. Such a world model must be quick to adapt to grid changes and needs to be able to cope with changes in marginal distributions.
2 Related Work
@@
The traditional and most widely-used approach for PSSE is the weighted least squares (WLS) method [1]. This algorithm minimizes the sum of the squared differences between the observed and estimated measurements, with each term being weighed inversely proportionally to the square of the measurement error standard deviation.
@@
However, the WLS algorithm is computationally intensive. Its time complexity is generally considered to be $\mathcal{O}(N^{3})$ in the number of buses $N$, assuming a dense system matrix. This is due to the need for matrix inversions and solving linear equations. This complexity can become a limitation for large-scale power systems with thousands of buses, leading to significant computational burden and time constraints, especially when real-time or near-real-time estimations are required. Additionally, WLS assumes that all error distributions are Gaussian, a condition that may not always hold true in practice.
@@
To overcome these limitations, an increasing number of publications instead use artificial neural networks (ANNs) for PSSE, a combination that is called NSE. ANNs may be able to perform the calculation faster than iterative solutions and achieve a higher solution quality simultaneously [13, 3]. However, like all machine learning (ML) methods, the performance of ANNs is contingent on the quality and quantity of the available training data. Therefore, NSE approaches are usually valid only for the grid they have been trained on. Once the topology or characteristics of nodes change, the ANN needs to be retrained. This is known as the problem of transfer learning (TL).
@@
The most logical way to overcome this limitation is to use models that incorporate information about the graph topology into their calculations. Such models are known under an umbrella term GNNs. Expectedly, recent years have seen a high volume of publications that propose utilizing GNNs for NSE in various ways. To name a few examples:
@@
  • [14] lays important groundwork in comparing different matrix representations of graphs within the graph convolutional network (GCN) model;
@@
  • [11] utilizes graph attention networks (GATs) with a different graph representation of the power grid;
@@
  • [9] explores the possibility of utilizing temporal correlations in the datasets using recurrent GCNs.
@@
However, to our knowledge, none of these research projects specifically considered the problem of ZSL in PSSE. The contribution of this work is in setting up multiple evaluation scenarios for ZSL and testing different configurations of GNNs in them.
3 Research question
@@
In the context of power grids, this is the difference between two use cases. In the homogeneous case, the power grid remains the same, but some connections between its nodes appear or disappear due to changes in switch states or elements going in and out of service. In the heterogeneous case, the model trained on one grid is used to make predictions about a completely different grid. Naturally, homogeneous TL capabilities are easier to achieve compared to heterogeneous.
This distinction becomes very important in production environments. Enabling the use of a model for a completely new grid naturally takes time, and training the model on that specific grid could be incorporated into this process without noticeably slowing it down. On the other hand, changes in grid topology due to switching can happen suddenly and unpredictably, and the model must adapt to them in real-time.
@@
A subset of TL is ZSL. This scenario excludes the possibility of fine-tuning the model on the new distribution and evaluates its performance directly after the transfer. In this project, we specifically focus on ZSL because it is more representative of real-life situations where a model must make predictions immediately after a topology change without access to any training data for fine-tuning.
Obviously, we cannot expect the model to perform equally well when transferring to a new distribution, be it a new topology or a different subset of observable buses. The goal is, rather, to ensure that the performance degradation is proportional to the magnitude of the change. In other words, the model should be resilient to distributional shifts.
Judging this by the usual error aggregation metrics proved challenging, especially when aggregation is performed over graph nodes, masking their differences. Analyzing the performance data without aggregation is no less challenging. Therefore, we developed a new metric that is hopefully more representative in the context of NSE. We define resilience as the number of buses in the grid for which the estimator predicts neither a false positive nor a false negative violation of the voltage band, divided by the total number of buses. The result is a number between zero and one, one being the ideal result. This metric is also discrete and scale-invariant, making it easier to comprehend. The downside is that it is easier to saturate, so different model configurations may exhibit the exact same performance.
Let $V_{n}^{min}:=0.9\cdot V_{n}^{rated}$ and $V_{n}^{max}:=1.1\cdot V_{n}^{rated}$ respectively be the lower and upper bound for the voltage band of node $n$. For simplicity, we will only consider the bounds of voltage magnitude and ignore the angle. Also, let $V_{n}$ and $\hat{V}_{n}$ be the ground truth and the predicted voltage magnitude at that node. Then the resilience $\mathcal{R}$ of a PSSE described by a vector of voltages $G$ can be formally defined as follows:
$\mathcal{R}(G)=\frac{1}{\|G\|}\sum_{n=1}^{\|G\|}\left[\begin{aligned} (V_{n}<V_{n}^{min}\land\hat{V}_{n}<V_{n}^{min})&\\ \lor(V_{n}^{min}\leq V_{n}\leq V_{n}^{max}\land&V_{n}^{min}\leq\hat{V}_{n}\leq V_{n}^{max})\\ \lor&(V_{n}>V_{n}^{max}\land\hat{V}_{n}>V_{n}^{max})\end{aligned}\right]$
This gives us the resilience value for each snapshot of the power grid, aggregated over all nodes. To then aggregate it over all snapshots, we take the 5th percentile of the results. This aggregation method focuses on representing the worst-case scenario while being robust to outliers.
The research question for this project is how to design a model for the PSSE problem that is resilient to changes in the data distribution, specifically:
@@
    To grid topology changes resulting from changing switch states.
@@
4 Methodology
4.1 Model selection
The general question of model selection for NSE was addressed by us previously in [4]. The main conclusion from that paper was the selection of GNNs as the most promising direction for further research. Now, we will perform a similar comparison study within the GNN family. We are comparing four models using the implementations provided by [6]:
@@ 4.1 Model selection
    Graph convolutional network (GCN)from [10]
@@
    Graph attention network (GAT)from [19]
@@
    Graph isomorphism network (GIN)from [21]
@@
    Graph sample and aggregate (GraphSAGE)from [7]
4.2 Graph representation of power systems
A successful application of GNN models naturally depends on how well the underlying data can be represented in the graph format. The most obvious first step is to represent buses in the grid as nodes of the graph and lines as its edges. In this project, we also represented transformers as edges without any additional parameters. For this to work, the voltage levels across the transformer must be normalized to avoid large voltage gradients.
@@ 4.2 Graph representation of power systems
4.3 Data preprocessing
GNN models are geometrically isodimensional, meaning that each output node must have a corresponding input node. This presents a problem for the PSSE use case where we lack input features for many if not most, input nodes. The question, therefore, is: How do we initialize the missing features in the input data?
The best solution seems to be the feature propagation algorithm from [15], which interpolates missing node-level features by solving a heat equation with known features as boundary conditions. This results in a smooth interpolation of features between known nodes and forms a good starting point for subsequent application of GNNs.
4.4 Datasets
The main dataset used in this project is the SimBench 1-MV-urban–1-sw, a 147-node, 10 kV medium voltage grid [12]. It is composed of a grid model and per-bus complex (active and reactive) power yearly time series with a resolution of one hour. To calculate the resulting grid state, we ran a power flow calculation using the energy system simulation SIMONA [8]. The resulting dataset comprises the base data and a year of complex voltage time series with a 15-minute temporal resolution. This dataset is hereafter called PQ.
Most grid branches in this model are of the open loop type, which means an open switch connects two separate branches. To simulate a realistic topology change, we made a line in one of the open loop branches inoperable, resembling a line fault, and closed the loop switch to resupply all nodes. Performing this operation on different branches resulted in multiple variations of the base grid topology. Afterward, we reran the simulation for each variation to obtain a topology change dataset, which is referenced hereafter as TC.
Unfortunately, the base dataset did not contain information about measurement devices. Therefore, we had to choose observable nodes randomly based on an observability level of 50%, which we assume is realistic for distribution grids.
An auxiliary dataset used in the heterogeneous ZSL experiments is based on the CIGRE medium voltage distribution network from Pandapower [17]. It is a much smaller grid with only 15 nodes. The voltage data for it is generated using the Midas simulation framework [2]. The shorthand name for this dataset is MV.
4.5 Use cases
@@
In the first use case corresponding to subquestion A, we train the model on the grid with a baseline level of observability and then linearly reduce it from the baseline level to zero at testing time. Of course, the model’s performance decreases along with this reduction. The shorthand name of this use case is observability degradation (OD).
The second use case corresponds to subquestion B and tests ZSL for homogeneous topology changes. In it, we split the TC dataset in a 50:50 ratio, train the model on the first part, and evaluate on the second. We also evaluate another model trained on the PQ dataset on the TC testing subset to see if the model needs to "see" the topology changes happening in order to be able to adapt to them at testing time, but our null hypothesis is that this is not the case. This scenario has the shorthand name TC1, and the former one where the model is trained on the TC dataset is called TC2.
The third use case corresponds to subquestion C and covers the heterogeneous ZSL scenario. Here, we transfer the model between PQ and MV datasets in both directions, i. e., training on one and then testing on another. The scenario where the model is trained on PQ and tested on MV has the shorthand name PQ2MV, and the other one has MV2PQ.
4.6 Experiment setup
Let us now establish the full hyperparameter space for the model in question. It consists of the following dimensions:
  • Model, as listed in Model selection. Categorical parameter with four values.
@@ 4.6 Experiment setup
  • Use of feature propagation (instead of initializing the missing features with zeros). Boolean parameter.
@@
Unfortunately, preliminary experiments have demonstrated that the hyperparameter space is not separable, meaning that a full grid search of the space is required. We performed this search for all model configurations and use cases and collected two metrics for each one: the resilience metric defined in Research question and mean squared error (MSE), which is used as the loss function for training the models. While resilience focuses on discrete voltage violations critical for practical applications, MSE is a continuous measure sensitive to small prediction errors. The results of this experiment are organized into an evaluation table where rows correspond to model configurations and columns consist of two groups:
@@
    Model configuration parameters
    1. (a)
      “model” is the name of the model from Model selection;
    2. (b)
@@
    5. (e)
      “#params” is not an independent parameter but a reference column containing the number of trainable parameters for the given model configuration.
  2. 2.
    Results of the form “usecase_metric”, where use case corresponds to one of the shorthand names from Use cases and metric is either “res” of “mse”, which stands for resilience and MSE, respectively.
The full table is available in our repository [^1] https://gitlab.com/transense/nse-tl-paper/blob/ECML/results.csv, and in the next section, we will use subsets of it as illustrations of results.
5 Evaluation
@@
    How is the performance on different tasks correlated?
@@
    How does model complexity affect performance?
@@
    How do the data augmentations proposed in Methodology affect performance?
  5. 5.
    Are resilience and MSE aligned?
5.1 Ranking model configurations
model layers fp adm od_mse od_res
GCN 1 True True 1.184553 1.000000
GCN 1 True False 1.195014 1.000000
GAT 6 False False 1.210447 1.000000
GAT 2 True False 1.233091 1.000000
GCN 2 True True 1.234646 1.000000
GAT 5 False True 1.265148 1.000000
GAT 2 True True 1.268001 1.000000
GAT 1 True False 1.279489 1.000000
GCN 2 True False 1.287740 1.000000
GAT 1 True True 1.293704 1.000000
Table 1: Best configurations for observability degradation
In this use case (table 1), models can saturate the resilience metric, so MSE becomes the main ranking factor. The winning models are GCN and GAT, and, curiously, mostly their shallow versions. It appears that shallow models perform better with data augmentation, especially feature propagation, while deeper models perform better without it. The reason for this could be the over-smoothing effect, which is a known weakness of deep GNNs. It occurs because as the number of layers increases, the node features tend to become more homogenized, leading to a loss of information. Feature propagation deliberately provides smooth data interpolation, further exacerbating this problem.
Best configurations for switching changes
Table 2: PQ →TC
model layers fp adm tc1_mse tc1_res
GAT 6 True False 1.654139 0.663366
GAT 4 True False 1.851676 0.663366
GAT 6 False False 1.979479 0.663366
GraphSAGE 3 True True 1.979515 0.663366
GAT 2 True False 1.983870 0.663366
GraphSAGE 2 True False 1.988077 0.663366
GAT 5 False True 1.991581 0.663366
GIN 2 True False 1.993644 0.663366
GraphSAGE 1 True True 1.995803 0.663366
GAT 3 True True 1.996154 0.663366
Table 3: TC →TC
model layers fp adm tc2_mse tc2_res
GAT 7 False False 1.853729 0.663366
GAT 6 False False 1.881083 0.663366
GraphSAGE 3 True True 1.899644 0.663366
GraphSAGE 2 True False 1.904953 0.663366
GCN 2 True False 1.919259 0.663366
GAT 2 True True 1.921153 0.663366
GraphSAGE 2 True True 1.939164 0.663366
GAT 2 True False 1.946193 0.663366
GIN 1 True True 1.958784 0.663366
GAT 4 False True 1.972278 0.663366
In both scenarios of the second use case (tables 3 and 3), all models, regardless of configuration, are approaching the same performance limit.
Best configurations for heterogeneous transfer
Table 4: PQ →MV
model layers fp adm pq2mv_mse pq2mv_res
GCN 1 False True 0.553909 1.000000
GCN 2 False True 0.867020 0.933333
GIN 1 False True 1.640098 0.933333
GAT 1 False True 1.487246 0.800000
GAT 2 True False 1.740887 0.933333
GraphSAGE 2 False True 1.743540 0.933333
GAT 1 False False 1.497050 0.800000
GCN 5 True True 1.782348 0.933333
GIN 2 False True 1.797467 0.933333
GIN 2 True True 1.808829 0.933333
Table 5: MV →PQ
model layers fp adm mv2pq_mse mv2pq_res
GraphSAGE 5 True True 0.014931 1.000000
GraphSAGE 3 True True 0.019819 1.000000
GraphSAGE 5 True False 0.021773 1.000000
GraphSAGE 7 True True 0.029667 1.000000
GraphSAGE 2 True True 0.030972 1.000000
GraphSAGE 4 False True 0.054912 1.000000
GAT 4 True True 0.055044 1.000000
GAT 2 False False 0.055671 1.000000
GraphSAGE 3 False False 0.055920 1.000000
GraphSAGE 2 False False 0.062780 1.000000
The third use case is the most surprising. While heterogeneous transfer should theoretically be a harder problem than the homogeneous one, here we see that the performance of models in this use case exceeds that in the previous one, and some models are even saturating the resilience metric. The second surprise is that a transfer from a smaller grid to a bigger one (MV with 15 buses to PQ with 147 buses) produces better results, and by a very large margin.
Before proposing theoretical explanations for these results, we must conduct further data analysis.
5.2 Correlation analysis
In this and the following subsections, we need to introduce a composite metric defined as the ratio of resilience to MSE. A larger value of this metric corresponds to better performance. The reasons for this are two. First, resilience by itself is not very conducive to correlation analysis since it is often saturated. Second, performing the same analysis for both metrics complicates the comprehension of results. We will study the Alignment of metrics separately later. Until then, if the metric used is not clarified, then it should be understood as the composite one.
We also introduce another use case as a baseline. In it, a model is trained on the first half of the PQ dataset and tested on the second half. It is meant to indicate how well the model can fit data without ZSL considerations. In figures it is labeled as “test_base”.

### v2

# references.bib

@@ Abstract
This paper addresses the challenge of neural state estimation in power distribution systems. We identified a research gap in the current state of the art, which lies in the inability of models to adapt to changes in the power grid, such as loss of sensors and branch switching, in a zero-shot fashion. Based on the literature, we identified graph neural networks as the most promising class of models for this use case. Our experiments confirm their robustness to some grid changes and also show that a deeper network does not always perform better. We propose data augmentations to improve performance and conduct a comprehensive grid search of different model configurations for common zero-shot learning scenarios.
Index Terms:
@@
neural state estimation; zero-shot learning; transfer learning; graph neural networks.
I Introduction
@@
Power System State Estimation (PSSE) is the task of inferring the “state” of an electrical power grid from real-time data collected by various sensors distributed across the system. The “state” in this context generally refers to the voltage magnitudes and phase angles at each bus in the grid.
@@
For many years, PSSE was mainly performed for the transmission grids using simplifying assumptions such as near-DC power flow and computational methods with poor scalability [wu1990power]. This is enabled by balanced operation with a relatively simple, predominantly linear topology of transmission grids, given their scale and structure.
@@
On the contrary, distribution grids, which transport electricity from substations to end consumers, present distinct challenges. Their unbalanced nature, radial or weakly meshed topology, high R/X ratios, and cost inefficiency to achieve sufficient sensor coverage complicate the state estimation process. Initially designed with transmission systems in mind, conventional methods often struggle to provide accurate state estimation in these more complex, dynamic, and less predictable distribution systems [wu1990power].
@@
However, with the proliferation of Distributed Energy Resources (DERs) and other complex consumers, grid operators are facing the necessity of performing PSSE for distribution grids. Additionally, §14a of the German Energy Industry Act effectively requires operators to develop transparency in distribution grids in order to align consumption with production from renewable energy sources, which requires PSSE.
In this paper, we begin by reviewing relevant prior work in Section II, followed by a formal statement of our research question in Section III. Section IV details the methodology, including model selection, data preprocessing, and experimental setup. We present and analyze our results in Section V and discuss their implications. Finally, Section VII summarizes our findings and suggests directions for future work.
II Related Work
@@
The traditional and most widely-used approach for PSSE is the Weighted Least Squares (WLS) method [wls]. This algorithm minimizes the sum of the squared differences between the observed and estimated measurements, with each term being weighed inversely proportionally to the square of the measurement error standard deviation.
@@
However, the WLS algorithm is computationally intensive. Its time complexity is generally considered to be $\mathcal{O}(N^{3})$ in the number of buses $N$, assuming a dense system matrix [wls]. This is due to the need for matrix inversions and solving linear equations. This complexity can become a limitation for large-scale power systems with thousands of buses, leading to significant computational burden and time constraints, especially when real-time or near-real-time estimations are required. Additionally, WLS assumes that all error distributions are Gaussian, a condition that may not always hold true in practice.
@@
To overcome these limitations, an increasing number of publications instead use Artificial Neural Networks (ANNs) for PSSE, a combination that is called Neural State Estimation (NSE). ANNs may be able to perform the calculation faster than iterative solutions and achieve a higher solution quality simultaneously [mestav2018state][balduin2020evaluating]. However, like all Machine Learning (ML) methods, the performance of ANNs is contingent on the quality and quantity of the available training data. Therefore, NSE approaches are usually valid only for the grid they have been trained on. Once the topology or characteristics of nodes change, the ANN needs to be retrained. This is known as the problem of Transfer Learning (TL).
@@
The most logical way to overcome this limitation is to use models that incorporate information about the graph topology into their calculations. Such models are known under an umbrella term Graph Neural Networks (GNNs). Expectedly, recent years have seen a high volume of publications that propose utilizing GNNs for NSE in various ways. To name a few examples:
@@
  • nse1 nse1 lays important groundwork in comparing different matrix representations of graphs within the Graph Convolutional Network (GCN) model;
@@
  • nse2 nse2 utilizes Graph Attention Networks (GATs) with a different graph representation of the power grid;
@@
  • nse3 nse3 explore the possibility of utilizing temporal correlations in the datasets using recurrent GCNs.
@@
However, to our knowledge, none of these research projects specifically considered the problem of Zero-Shot Learning (ZSL) in PSSE. The contribution of this work is in setting up multiple evaluation scenarios for ZSL and testing different configurations of GNNs in them.
III Research question
@@
In the context of power grids, this is the difference between two use cases. In the homogeneous case, the power grid remains the same, but some connections between its nodes appear or disappear due to changes in switch states or elements going in and out of service. In the heterogeneous case, the model trained on one grid is used to make predictions about a completely different grid [Yang2021Power-grid].
This distinction becomes very important in production environments. Integrating a model into the control system of a real grid naturally takes time, and training the model on that specific grid could be incorporated into this process without noticeably slowing it down. On the other hand, changes in grid topology due to switching can happen suddenly and unpredictably, and the model must adapt to them in real-time.
@@
A subset of TL is Zero-Shot Learning (ZSL). This scenario excludes the possibility of fine-tuning the model on the new distribution and evaluates its performance directly after the transfer. In this project, we specifically focus on ZSL because it is more representative of real-life situations where a model must make predictions immediately after a topology change without access to any training data for fine-tuning. In other words, the model should be robust to distributional shifts.
Of course, in practice, a model can be fine-tuned to provide the best performance for the new topology. Still, until this process is complete, the previous version of the model has to substitute for it and provide good enough estimations, even if they are of lower quality.
The research question for this paper is which existing models in application to the PSSE problem are robust to changes in the data distribution, specifically:
@@
    To grid topology changes resulting from changing switch states;
@@
IV Methodology
IV-A Model selection
The general question of model selection for NSE was addressed by us previously in [modelcomparison]. The main conclusion from that paper was the selection of GNNs as the most promising direction for further research. Now, we will perform a similar comparison study within the GNN family. We are comparing four models using the implementations provided by PyTorch Geometric framework [pyg]:
@@ IV-A Model selection
    Graph Convolutional Network (GCN)as proposed in [gcn]
@@
    Graph Attention Network (GAT)as proposed in [gat]
@@
    Graph Isomorphism Network (GIN)as proposed in [gin]
@@
    Graph Sample and Aggregate (GraphSAGE)as proposed in [sage]
IV-B Graph representation of power systems
A successful application of GNN models naturally depends on how well the underlying data can be represented in the graph format. The first step is to represent buses in the grid as nodes of the graph and lines as its edges. In this project, we also represented transformers as edges without any additional parameters. For this to work, the voltage levels across the transformer must be normalized to avoid large voltage gradients.
@@ IV-B Graph representation of power systems
IV-C Data preprocessing
GNN models are geometrically isodimensional, meaning that each output node must have a corresponding input node. This presents a problem for the PSSE use case, where we lack input features for many, if not most, input nodes. The question, therefore, is: How do we initialize the missing features in the input data?
The solution we chose is the feature propagation algorithm from [rossi2021fp], which interpolates missing node-level features by solving a heat equation with known features as boundary conditions. This results in a smooth interpolation of features between known nodes and forms a starting point for the subsequent application of GNNs.
IV-D Datasets
Figure 1: A visualization of the SimBench 1-MV-urban–1-sw grid
The main dataset used in this project is the SimBench 1-MV-urban–1-sw, a 147-node, 10 kV medium voltage grid [meineckeSimBenchBenchmarkDataset2020] depicted in Figure 1. It is composed of a grid model and a per-bus complex (active and reactive) power yearly time series. To calculate the resulting grid state, we performed a power flow calculation using SIMONA energy system simulation [hiryAgentbasedDiscreteEventSimulation2022]. The resulting dataset comprises the base data and a year of complex voltage time series with a 15-minute temporal resolution. This dataset is hereafter called PQ.
Most grid branches in this model are of the open loop type, which means an open switch (depicted as a square) connects two separate branches. To simulate a realistic topology change, we made a line in one of the open loop branches inoperable, resembling a line fault, and closed the loop switch to resupply all nodes. Performing this operation on different branches resulted in multiple variations of the base grid topology. Afterward, we reran the simulation for each variation to obtain a topology change dataset, which is referenced hereafter as TC.
Unfortunately, the base dataset did not contain information about measurement devices. Therefore, we had to choose observable nodes randomly based on an observability level of 50%, which we assume is realistic for distribution grids. This means that the state estimator has access to true voltage values for half of the grid buses.
An auxiliary dataset used in the heterogeneous ZSL experiments is based on the CIGRE medium voltage distribution network from Pandapower [pandapower]. It is a much smaller grid with only 15 nodes, which allows us to study how the complexity of the grids affects the performance of ZSL. The voltage data for it is generated using the Midas simulation framework [midas]. The shorthand name for this dataset is MV.
IV-E Use cases
@@
In the first use case corresponding to subquestion A, we train the model on the grid with a baseline level of observability and then linearly reduce it from the baseline level to zero at testing time. Of course, the model performance decreases along with this reduction. The shorthand name of this use case is observability degradation (OD).
The second use case corresponds to subquestion B and tests ZSL for homogeneous topology changes. In it, we split the TC dataset in a 50:50 ratio, train the model on the first part, and evaluate on the second. We also evaluate another model trained on the PQ dataset on the TC testing subset to see if the model needs to observe the topology changes happening in order to be able to adapt to them at testing time, but our null hypothesis is that this is not the case. This scenario has the shorthand name TC1, and the former, where the model is trained on the TC dataset, is called TC2.
The third use case corresponds to subquestion C and covers the heterogeneous ZSL scenario. Here, we transfer the model between the PQ and MV datasets in both directions, that is, training on one and then testing on another. The scenario in which the model is trained on PQ and tested on MV has the shorthand name PQ2MV, and the other has MV2PQ.
IV-F Experiment setup
Let us now establish the full hyperparameter space for the models in question. It consists of the following dimensions:
  • Model, as listed in the Model selection subsection. Categorical parameter with four values.
@@ IV-F Experiment setup
  • Use of feature propagation (as opposed to initializing the missing features with zeros). Boolean parameter.
@@
Unfortunately, preliminary experiments have demonstrated that the hyperparameter space is not separable, meaning that a full grid search of the space is required. We performed this search for all model configurations and use cases and collected the Mean Squared Error (MSE) metric for each one. The results of this experiment are organized into an evaluation table where rows correspond to model configurations and columns are the following:
@@
    “model” is the name of the model;
  2. 2.
@@
  5. 5.
    “mse” is the value of MSE for the configuration defined by the above parameters.
The full table is available in our repository in the “results.csv” file, and in the next section, we will use subsets of it as illustrations of results.
V Evaluation
@@
    How does model complexity affect performance?
@@
    How do the data augmentations proposed in the Methodology section affect performance?
@@
    How is performance on different tasks correlated?
The answers to these questions will then be used to answer the main research questions from Section III.
V-A Ranking model configurations
To interpret the numerical results listed in this section, it is useful to keep in mind the baseline value of MSE obtained by evaluating the models trained on the first half of the PQ dataset on the second half of the same dataset and taking the best result. This value is $0.32$. We can then broadly say that ZSL is possible in scenarios where the value of MSE after the topology change does not significantly exceed it.
TABLE I: BEST CONFIGURATIONS FOR OBSERVABILITY DEGRADATION (OD)
model layers fp adm mse
GraphSAGE 3 True False 0.86
GraphSAGE 2 True True 0.87
GraphSAGE 3 True True 0.87
GraphSAGE 2 True False 0.88
GCN 3 True False 0.90
The winning model for the first use case is Graph Sample and Aggregate (GraphSAGE) utilizing feature propagation. It also appears from Table I that there is a sweet spot in model complexity of 2-3 layers.
TABLE II: BEST CONFIGURATIONS FOR SWITCHING CHANGES (TC1)
model layers fp adm mse
GraphSAGE 3 True False 0.31
GraphSAGE 1 True True 0.33
GAT 1 True False 0.33
GraphSAGE 2 True False 0.34
GraphSAGE 3 True True 0.34
TABLE III: BEST CONFIGURATIONS FOR SWITCHING CHANGES (TC2)
model layers fp adm mse
GraphSAGE 1 True False 0.32
GraphSAGE 1 True True 0.32
GAT 3 True False 0.32
GAT 1 True True 0.32
GAT 1 True False 0.32
In both scenarios of the second use case (Tables II and III), the model rating is similar, which suggests that the tasks themselves are similar as well. The winning models are GraphSAGE and GAT, also with the help of feature propagation and, curiously, in their shallowest versions, with single-layer models showing some of the best results. The main observation, however, is that the MSE values are identical to the baseline, meaning that homogeneous ZSL works very well.
TABLE IV: BEST CONFIGURATIONS FOR HETEROGENEOUS TRANSFER (PQ2MV)
model layers fp adm mse
GCN 1 False True 0.30
GCN 2 False True 0.87
GCN 4 False True 0.97
GAT 2 False False 1.06
GraphSAGE 2 False True 1.26
TABLE V: BEST CONFIGURATIONS FOR HETEROGENEOUS TRANSFER (MV2PQ)
model layers fp adm mse
GAT 6 True False 0.62
GAT 2 True False 0.64
GCN 2 True True 0.67
GIN 1 True True 0.69
GAT 2 False True 0.78
In the third use case, we see a significant difference between the two scenarios. In the first scenario (Table IV), where the model is transferred from a larger to a smaller grid, the best-performing by a large margin is a single-layer GCN, which is the simplest of all the compared models. A possible explanation is that there are only a few correlations that are reusable between grids, which the simple model can capture. Any more complex model picks up too many correlations that are specific to the grid it was trained on and then misapplies them. In the second scenario (Table V), the results are mixed between complex and simple models, and we cannot come up with a sound theoretical interpretation of this result.
V-B Impact of model complexity
Figure 2: Performance as a function of model complexity
In Figure 2, we plot the performance of models against their trainable parameter counts. We also plot the average for all models of a given complexity: the “Mean” line on the graph.
Here, we can see a break point at about 84 parameters or 8 layers, starting from which the performance of models becomes much more consistent between use cases and configurations. The explanation for this effect is the over-smoothing phenomenon described in [oversmoothing]. In short, GNN layers of all architectures tend to act as low-pass filters, which effectively averages the output values over multiple iterations. Eventually, the model converges to an output where the values at all nodes of the graph are identical.
Since we are not interested in over-smoothed results, we can drop the model configurations that output them from further analysis. Therefore, the following sections will use the results table truncated to a maximum of 7 layers.
V-C Impact of data augmentations
TABLE VI: IMPACT OF DATA AUGMENTATIONS
fp adm od tc1 tc2 pq2mv mv2pq
True True 1.33 1.17 1.13 2.19 1.80
True False 1.25 1.03 1.06 1.99 2.55
False True 1.51 1.24 1.25 1.93 2.00
False False 1.49 1.23 1.10 1.98 2.09
In Table VI, we average the performance values across all model configurations, leaving only the data augmentations as parameters. The results are unsurprising and follow the observations we made previously. Homogeneous scenarios benefit from feature propagation but are held back by admittance weights. In the heterogeneous scenarios, we once again see a split where MV2PQ benefits from admittance weights and PQ2MV does not. On average, the impact of the augmentations is not very significant.
V-D Correlation analysis